The standard deviation of the market-index portfolio is 40%. Stock A has a beta of 1.75 and a residual standard deviation of 50%.

a. Calculate the total variance for an increase of 0.10 in its beta. (Do not round intermediate calculations. Round your answer to the nearest whole number.)

b. Calculate the total variance for an increase of 5.46% in its residual standard deviation. (Do not round intermediate calculations. Round your answer to the nearest whole number.)
Answer #1

Given ,β= 1.75 and σ(e) = 0.50 (50%)

Total variance normally = Systematic variance + Residual variance = β^2 * Var(rM) + Var(e)

variance = (1.75)^2 * (0.40)^2 + (0.50)^2

= (1.75 * 1.75) * (0.40 * 0.40 ) + (0.50 * 0.50 )

= 0.49 + 0.25

= 0.79

(a) When beta is increased by 0.10, the new data will be  

β= 1.85 and σ(e) = 0.50

The total variance = (1.85)^2 * (0.40)^2 + (0.50)^2

= (1.85 * 1.85) * (0.40 * 0.40 ) + (0.50 * 0.50 )

= 0.5476 + 0.25

= 0.7976

(b) When the residual standard deviation is increased by 5.46%

  β= 1.75 and σ(e) = 0.5546 (0.50+ 0.0546 increase in residual stnd. deviation)

Total variance = (1.75)^2 * (0.40)^2 + (0.5546)^2

= (1.75 * 1.75) * (0.40 * 0.40 ) + (0.5546 * 0.5546 )

= 0.49 + 0.30758

= 0.79758

= 0.7976 (approximatly)

Both the section (a) and (b) have same impact .The total variance will increase from 0.79 to 0.7976.